OpenStudy (anonymous):

An employer must hire 3 programmers from among 6 applicants and 1 product manager from among 4 applicants. What is a total number of possible ways in which he can make his selection?

OpenStudy (anonymous):

use combinations C(6,3)*C(4,1) =20*4=80

OpenStudy (anonymous):

thanks